ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 128, 2002

Aaronson, J., H. Nakada, O. Sarig and R. Solomyak Invariant measures and asymptotics for some skew products

93

Ballico, E. Holomorphic vector bundles on C^{2}\{0}  197

Boland, J. and C. Connell Minimal entropy rigidity for foliations of compact spaces

221

Connell, C. See J. Boland

De Azevedo, S. S. See P. Koshlukov

Dijkstra, J. J. and J. van Mill Topological equivalence of discontinuous norms

177

Domingo-Juan, M. C., X. Gual and V. Miquel Pappus type theorems for hypersurfaces in a space form

205

Gasparis, I. On contractively complemented subspaces of separable L_{1}-preduals

77

Grunenfelder, L., T. Košir, M. Omladič and H. Radjavi Maximal Jordan algebras of matrices with bounded number of eigenvalues

53

Gual, X. See M. C. Domingo-Juan

Klimenko, E. and N. Kopteva Discreteness criteria for RP groups 247

Kopteva, N. See E. Klimenko

Koshlukov, P. and S. S. de Azevedo Graded identities for T -prime algebras over fields of positive characteristic

157

Košir, T. See L. Grunenfelder

Leinen, F. and O. Puglisi Confined subgroups in periodic simple finitary linear groups

285

MacCluer, B. and K.Saxe Spectra of composition operators on the Bloch and Bergman spaces

325

Miquel, V. See M. C. Domingo-Juan

Nakada, H. See J. Aaronson

Neunhäuserer, J. Number theoretical peculiarities in the dimension theory of dynamical systems

267

Omladič, M. See L. Grunenfelder

Puglisi, O. See F. Leinen Radjavi, H. See L. Grunenfelder Rozenblat, B. V. Elementary theories of completely simple semi-

groups

355

Sarig, O. See J. Aaronson Saxe, K. See B. MacCluer Solomyak, R. See J. Aaronson

Steinmetz, N. Value distribution of the Painlevé transcendents  29
Sun, Z.-W. On the sum ∑_{k≡r (mod m)}\binom{n}{k} and related congruences  135
Väisälä, J. Isometric approximation property in euclidean spaces  1

Van Mill, J. See J. J. Dijkstra